OpenStudy (anonymous):

3log(2x+2)=x Round to the nearest hundredth. If there is more than one solution, separate them with commas.

OpenStudy (isaiah.feynman):

Remember your log rules? @gabriel72alv

OpenStudy (anonymous):

I haven't done log in a while? /:

OpenStudy (anonymous):

please check your question. \[\log(2x+2)=\frac{ x }{ 3 },10^{\frac{ x }{ 3 }}=2x+2\]